# Adaptive Extraction Network for Multivariate Long Sequence Time-Series Forecasting

### v1

@@ 3.4 Dynamic prediction
Model forecasting. In the downstream prediction tasks of ATVCNet, we employ a simple feedforward neural network (FFN) architecture. To address potential underfitting issues in regression tasks and improve the computational efficiency of the model, we utilize the Moore-Penrose pseudo-inverse matrix [40] to compute the parameters of the model network quickly. The mathematical expression for the model prediction task is as follows:
@@
Model dynamic updates. As time progresses, the data distribution in real-world applications may change, resulting in a significant deterioration of the model’s predictive performance. To address this issue, we extend the model training process to a dynamic update stage. We assume that when the model’s predictive performance decreases by 5%, we adopt newly collected samples to retrain the ATVCNet model. It is noteworthy that during this process, we maintain a fixed training set size, which means that when introducing new data, an equal amount of the earliest historical data is removed. This strategy ensures the speed of model training and reduces the required time.
@@
1:  Phase I. Training the proposed ATVCNet model.
@@
8:  Input the obtained new samples $x_{j}$ and training data set $X$ into the trained ATVCNet model. If the MSE value exceeds the threshold, update $\beta$.
@@
Algorithm 1 Dynamic prediction.
@@ 4 Experiments
In this section, we delve into investigating the effectiveness and efficiency of the ATVCNet network. Through rigorous experimentation across eight authentic datasets, employing fourteen SOTA models, we aim to address the following research inquiries:
  • RQ1 (Effectiveness): Can ATVCNet outshine the present SOTA baseline models when applied to real datasets (Sections 4.4 - 4.7)?
  • RQ2 (Efficiency): Does ATVCNet exhibit superior performance in terms of resource utilization compared to the current baseline models (Sections 4.8)?
  • RQ3 (Ablation): To what extent do the distinct components of ATVCNet influence its overall performance in time series forecasting tasks (Sections 4.9)?
@@ 4.1 Datasets
We extensively evaluated the proposed ATVCNet model on eight benchmark datasets spanning four real-world fields: energy, finance, medical, and traffic. Table 1 summarizes the key features of these datasets.
  • ETT [^1] ETT dataset was acquired at https:// github.com/zhouhaoyi/ETDataset: The dataset records continuous operation data of power resources over a period of two years, including seven indicators such as oil temperature and load. ETTm1 and ETTm2 denote measurements taken every 15 minutes, while ETTh1 and ETTh2 represent hourly recordings.
@@
Table 1: The key features of the eight time series datasets.
Datasets Features Timesteps Granularity Information
ETTh1/h2 7 17420 1 hour Electricity
@@ 4.2 Baselines
We selected 14 representative SOTA models that have demonstrated outstanding performance in the field of MLSTF as benchmarks:
(1) MLP-based: TimeMixer [31] conducts long-term forecasting by decoupling historical information of complex patterns in multiscale time series. TSMixer [12] combines Patch and MLP, enabling it to extract both temporal correlations and inter-channel correlations. DLinear [10] decomposes time series and utilizes linear network layers for modeling to achieve prediction.
(2) Transformer-based architectures: iTransforemr [8] effectively captures multivariate correlations by independently embedding each variable in the time series as variable sub-tokens. PatchTST [7] employs a channel independence strategy, specifically designed to extract correlations of each channel in multivariate data. Crossformer [9] captures temporal and cross-dimensional dependencies in multivariate time series data by first transforming it into a 2D array and then employing two-stage attention layers. ETSformer [24] and FEDformer [25] both decompose time series data to extract complex patterns. Non-Stationary [41] incorporating sequence stabilization and de-stabilization attention mechanisms to handle non-stationary information.
(3) CNN-based model: ConvTimeNet [13] adaptively segments time series into patches and integrates deepwise convolution and pointwise convolution operations to capture global sequence dependencies and cross-variable interactions. MICN [14] utilizes multi-scale convolution to extract both global and local contextual relationships.
(4) 2D-variation-based: TimesNet [42] converts the 1D temporal dimension into 2D to extract both intra-segment dependencies and dependencies between adjacent segments in time.
(5) Semantic-based model: TS2Vec [43] employs hierarchical contrastive methods on both instance and time dimensions to capture multiscale contextual relationships.
(6) RNN-based model: LSSL [20] simulates basic state space models, utilizing novel parameterization methods to model structured state space sequences, in order to extract longer dependencies.
@@ 4.3 Implementation details
Our approach is trained with an initial learning rate of $1\times 10^{-3}$, the convolution dilated rate $d$ for multi-resolution dilated convolution was set to $\left\{1,2,5\right\}$. The input sequence length for the ILI dataset is $L=36$, with prediction sequence lengths of $\left\{24,36,48,60\right\}$. For other datasets, the input sequence length is $L=96$, with prediction sequence lengths of $\left\{96,192,336,720\right\}$. To verify the robustness of our results, we trained the ATVCNet model with three different random seeds and calculated the mean squared error (MSE) and mean absolute error (MAE) scores using each selected seed. The average results are recorded as shown in Table 2. Additionally, the efficiency of the model is assessed based on floating point operations (FLOPs), Parameters, training time, and Inference time. Experiments are implemented using PyTorch on a single NVIDIA GeForce RTX 3090 24GB GPU, Intel(R) Core(TM) i7-10700k CPU and 32GB RAM.
4.4 Multivariate results and analysis
@@
In the field of multivariate prediction, the ATVCNet framework demonstrates excellent performance on almost all datasets. Through analysis of the result data in Table 2, the following conclusions can be drawn:
@@
  • The ATVCNet model significantly improves inference performance on nearly all datasets. As the forecasting horizon increases, the prediction error of ATVCNet shows a gradually stable upward trend, indicating its ability to maintain high long-term robustness in practical applications.
@@
  • Compared to the current SOTA forecasting model ConvTimeNet, ATVCNet exhibits notable improvements across various datasets. Specifically, ATVCNet gives 47.9% $\left(0.319\to 0.166\right)$ MSE reduction in ETTm1, 36.9% $\left(0.211\to 0.133\right)$ in ETTm2, 42% $\left(0.395\to 0.229\right)$ in ETTh1, 58% $\left(0.436\to 0.183\right)$ in ETTh2, 32.1% $\left(0.202\to 0.137\right)$ in Electricity, 39.8% $\left(0.402\to 0.242\right)$ in Traffic, 65.7% $\left(0.347\to 0.119\right)$ in exchange and 82.9% $\left(1.866\to 0.319\right)$ in ILI. Overall, ATVCNet yields a 63.4% averaged MSE reduction among above settings. Notably, our model achieved great success on the electricity dataset (321 variables) and the traffic dataset (862 variables), indicating that, compared to the ConvTimeNet model which uses depthwise separable convolutions to extract correlations between different variables, the inter-variable dependency extraction at different resolution scales by ATVCNet is more beneficial for MLSTF.
@@
  • ATVCNet outperforms both the iTransformer and TSMixer models significantly. Compared to iTransformer, ATVCNet reduces the MSE across all datasets by 66.8% $\left(0.576\to 0.191\right)$. Similarly, compared to TSMixer, ATVCNet reduces the MSE across all datasets by 49.6% $\left(0.379\to 0.191\right)$. iTransformer and TSMixer both emphasize feature extraction between different variables in time series. While they perform better than models such as PatchTST, which focuses solely on extracting temporal information, they may be affected by strong correlations between different variables, leading to information redundancy and decreased prediction accuracy. ATVCNet uses deformable convolution to extract the correlation between different variables in the time series and uses a flexible warp network to focus on important features, remove redundant features, and improve the prediction performance of the model.
@@
  • The ATVCNet model achieves significantly better results than the SOTA MLP-based models. ATVCNet has an average MSE decrease of 67.5% $\left(0.587\to 0.191\right)$ on all datasets in TimeMixer and 49.6% $\left(0.379\to 0.191\right)$ in TSMixer. We analyzed that this may be because MLP-based models need to extract more feature information from complex time series to fit future trends. Thanks to the comprehensive feature extraction capability of the model, ATVCNet can not only extract and utilize local contextual information and global correlation information of time series but also extract information between variables at different resolutions as a complement to MLSTF, which improves the prediction accuracy.
@@
In summary, the excellent performance of ATVCNet in handling long time series may be attributed to its comprehensive consideration of local information, global information, and interdependencies between different variables in the time series, aspects that many current MLSTF models do not take into account. This allows ATVCNet to excel in handling various real-world time series prediction tasks.
@@
Table 2: Multivariate results with different prediction lengths. The best results in bold numbers. Avg represents the average value across the four prediction lengths.

### v2

@@ 3.4 Dynamic prediction
Model prediction. In the downstream prediction tasks of ACNet, we employ a simple feedforward neural network (FFN) architecture. To address potential underfitting issues in regression tasks and improve the computational efficiency of the model, we utilize the Moore-Penrose pseudo-inverse matrix [38] to compute the parameters of the model network quickly. The mathematical expression for the model prediction task is as follows:
@@
Model dynamic updates. As time progresses, the data distribution in real-world applications may change, resulting in a significant deterioration of the model’s predictive performance. To address this issue, we extend the model training process to a dynamic update stage. We assume that when the model’s predictive performance decreases by 5%, we adopt newly collected samples to retrain the ACNet model. It is noteworthy that during this process, we maintain a fixed training set size, which means that when introducing new data, an equal amount of the earliest historical data is removed. This strategy ensures the speed of model training and reduces the required time.
@@
1:  Phase I. Training the proposed ACNet model.
@@
8:  Input the obtained new samples $x_{j}$ and training data set $X$ into the trained ACNet model. If the MSE value exceeds the threshold, update $\beta$.
@@
Algorithm 1 : Dynamic prediction.
@@ 4 Experiments
In this section, we will delve into the effectiveness and efficiency of the ACNet network. Through rigorous experiments on long-term and short-term forecasting using twelve real-world datasets and comparing against fifteen SOTA models, our goal is to address the following research questions:
  • RQ1 (Effectiveness): Can ACNet outshine the present SOTA baseline models when applied to real datasets (Sections 4.4 - 4.8)?
  • RQ2 (Efficiency): Does ACNet exhibit superior performance in terms of resource utilization compared to the current baseline models (Sections 4.9)?
  • RQ3 (Ablation): To what extent do the distinct components of ACNet influence its overall performance in time series forecasting tasks (Sections 4.10)?
@@ 4.1 Datasets
We extensively evaluated the proposed ACNet model on twelve benchmark datasets spanning four real-world fields: energy, finance, medical, and traffic. Table 1 summarizes the key features of these datasets.
  • ETT [^1] ETT dataset was acquired at https:// github.com/zhouhaoyi/ETDataset: The dataset records continuous operation data of power resources over a period of two years, including seven indicators such as oil temperature and load.
@@
  • PEMS[39]: This dataset records traffic network data collected every 5 minutes on California highways, including metrics such as traffic flow, speed, and occupancy.
Table 1: The key features of the twelve time series datasets.
Tasks Datasets Dim Timesteps Granularity Information
Long-term Forecasting ETTh1/h2 7 17420 1 hour Electricity
@@
Short-term Forecasting PEMS03 358 26208 5 min Traffic
PEMS04 307 16992 5 min Traffic
PEMS07 883 28224 5 min Traffic
PEMS08 170 17856 5 min Traffic
@@ 4.2 Baselines
We selected 15 representative SOTA models that have demonstrated outstanding performance in the field of TSF as benchmarks: Patch-based TSMixer [11], PatchTST [7], and PDF [40]; MLP-based TimeMixer [26], FITS [41], and DLinear [10]; Transformer-based architectures iTransformer [8], Crossformer [9], ETSformer [22], FEDformer [23], and Non-Stationary Transformer [42]; and CNN-based models ConvTimeNet [13], TSLANet [20], MICN [14], and TimesNet [43].
@@ 4.3 Implementation details
Our approach is trained with an initial learning rate of $1\times 10^{-3}$, the convolution dilated rate $d$ for multi-resolution dilated convolution was set to $\left\{1,2,5\right\}$. In long-term forecasting, the input sequence length for the ILI dataset is $L=36$, with prediction sequence lengths of $\left\{24,36,48,60\right\}$. For other datasets, the input sequence length is $L=96$, with prediction sequence lengths of $\left\{96,192,336,720\right\}$. In short-term forecasting, the input sequence length for the PEMS datasets is $L=96$, with a prediction sequence length of $12$. To verify the robustness of our results, we conducted long-term forecasting by training the ACNet model with three different random seeds. For each seed, we calculated the mean squared error (MSE) and mean absolute error (MAE). The average results are presented in Table 2. For short-term forecasting, we employed mean absolute error (MAE), mean absolute percentage error (MAPE), and root mean squared error (RMSE) as evaluation metrics. Additionally, we assessed the model’s efficiency based on floating point operations (FLOPs), number of parameters, training time, and inference time. Experiments are implemented using PyTorch on a single NVIDIA GeForce RTX 3090 24GB GPU, Intel(R) Core(TM) i7-10700k CPU and 32GB RAM.
4.4 Main results and analysis
@@
Long-term forecasting. In the field of long-term forecasting, the ACNet framework demonstrates excellent performance on almost all datasets. Through analysis of the result data in Table 2, the following conclusions can be drawn:
@@
  • The ACNet model significantly improves inference performance on nearly all datasets. As the forecasting horizon increases, the prediction error of ACNet shows a gradually stable upward trend, indicating its ability to maintain high long-term robustness in practical applications.
@@
  • Compared to the current SOTA forecasting model ConvTimeNet, ACNet exhibits notable improvements across various datasets. Specifically, ACNet gives 47.9% $\left(0.319\to 0.166\right)$ MSE reduction in ETTm1, 36.9% $\left(0.211\to 0.133\right)$ in ETTm2, 42% $\left(0.395\to 0.229\right)$ in ETTh1, 58% $\left(0.436\to 0.183\right)$ in ETTh2, 32.1% $\left(0.202\to\right.\\ \left.0.137\right)$ in Electricity, 39.8% $\left(0.402\to 0.242\right)$ in Traffic, 65.7% $\left(0.347\to 0.119\right)$ in exchange and 82.9% $\left(1.866\to 0.319\right)$ in ILI. Overall, ACNet yields a 63.4% averaged MSE reduction among above settings.
    Notably, ACNet has achieved great success on the Exchange and ILI datasets, which contain a large number of nonlinear features. This demonstrates the effectiveness of the ACNet model in handling nonlinear features. Notably, ACNet achieved significant success on the Exchange and ILI datasets, which contain a large number of nonlinear features. Compared to the ConvTimeNet model, which uses depthwise separable convolutions to extract time series features, ACNet’s extraction of nonlinear features and inter-variable correlations at different resolution scales proves to be more advantageous for TSF.
@@
  • ACNet outperforms both the iTransformer and TSMixer models significantly. Compared to iTransformer, ACNet reduces the MSE across all datasets by 66.8% $\left(0.576\to 0.191\right)$. Similarly, compared to TSMixer, ACNet reduces the MSE across all datasets by 49.6% $\left(0.379\to 0.191\right)$. iTransformer and TSMixer both emphasize feature extraction between different variables in time series. While they perform better than models such as PatchTST, which focuses solely on extracting temporal information, they may be affected by strong correlations between different variables, leading to information redundancy and decreased prediction accuracy. ACNet uses deformable convolution to extract the correlation between different variables in the time series and uses a flexible warp network to focus on important features, remove redundant features, and improve the prediction performance of the model.
@@
  • The ACNet model achieves significantly better results than the SOTA MLP-based models. ACNet has an average MSE decrease of 67.5% $\left(0.587\to 0.191\right)$ on all datasets in TimeMixer and 49.6% $\left(0.379\to 0.191\right)$ in TSMixer. We analyzed that this may be because MLP-based models need to extract more feature information from complex time series to fit future trends. Due to the comprehensive feature extraction capability of the model, ACNet can not only extract and utilize local contextual information and global correlation information of time series but also extract information between variables at different resolutions as a complement to TSF, which improves the prediction accuracy.
@@
In summary, the excellent performance of ACNet in handling long time series may be attributed to its comprehensive consideration of local information, global information, and interdependencies between different variables in the time series, aspects that many current TSF models do not take into account. This allows ACNet to excel in handling various real-world time series prediction tasks.
@@
Table 2: Long-term multivariate forecasting results with different prediction lengths. The best results are in bold numbers. Avg represents the average value across the four prediction lengths.